### v2

# Unifying Search and Recommendation: A Generative Paradigm Inspired by Information Theory

Abstract.
Recommender systems and search engines serve as foundational elements of online platforms, with the former delivering information proactively and the latter enabling users to seek information actively. Unifying both tasks in a shared model is promising since it can enhance user modeling and item understanding. Previous approaches mainly follow a discriminative paradigm, utilizing shared encoders to process input features and task-specific heads to perform each task. However, this paradigm encounters two key challenges: gradient conflict and manual design complexity. From the information theory perspective, these challenges potentially both stem from the same issue — low mutual information between the input features and task-specific outputs during the optimization process.
To tackle these issues, we propose GenSR, a novel generative paradigm for unifying search and recommendation (S&R), which leverages task-specific prompts to partition the model’s parameter space into subspaces, thereby enhancing mutual information. To construct effective subspaces for each task, GenSR first prepares informative representations for each subspace and then optimizes both subspaces in one unified model. Specifically, GenSR consists of two main modules: (1) Dual Representation Learning, which independently models collaborative and semantic historical information to derive expressive item representations; and (2) S&R Task Unifying, which utilizes contrastive learning together with instruction tuning to generate task-specific outputs effectively. Extensive experiments on two public datasets show GenSR outperforms state-of-the-art methods across S&R tasks. Our work introduces a new generative paradigm compared with previous discriminative methods and establishes its superiority from the mutual information perspective.
1. Introduction
Recommender systems and search engines have become indispensable components of modern online service platforms, including e-commerce websites and social media networks (Zhang et al., 2024; Wu et al., 2024). Usually, search and recommendation (S&R) tasks are trained on separate models (Bhattacharya et al., 2024). However, unifying these two tasks within a shared model offers significant advantages (Yao et al., 2021; Si et al., 2023): First, joint modeling of user behaviors across both S&R provides a more comprehensive understanding of user preferences. Second, leveraging interactions from both task domains enriches item-side interactions, enhancing the understanding and delivery of items. Therefore, it is promising to develop a unified model for S&R.
Figure 1. (a) and (b) illustrate the discriminative and generative paradigms for unifying S&R, respectively, highlighting their differences in parameter space. While the discriminative paradigm results in low mutual information by optimizing S&R tasks within a shared parameter space, the generative paradigm enhances mutual information by partitioning the space into distinct subspaces through prompts.
Existing efforts to unify S&R mostly follow a discriminative paradigm, which utilizes shared encoders to process input features (e.g., user interaction history) and task-specific heads to predict whether a user will click on an item in S&R tasks, as illustrated in Figure 1(a) (Shi et al., 2024; Xie et al., 2024). However, such paradigm has two primary drawbacks: 1) Gradient conflict: The optimization of shared parameters may introduce gradient conflict, thereby limiting the model’s performance (Shi et al., 2023; He et al., 2024). 2) Manual design complexity: Determining the extent of parameter sharing and designing task-specific heads requires significant manual effort, increasing complexity and limiting extensibility (Rahimian et al., 2023).
Inspired by information theory, we argue that these drawbacks potentially both stem from the same underlying issue — low mutual information between the model’s input features and task-specific output during optimization. This low mutual information arises from constraining both tasks within a shared parameter space in the shared encoder, as depicted in the parameter space analysis in Figure 1(a). This constraint limits the representation capability of the shared encoders, resulting in gradient conflict. Thus, it necessitates the manual design of task-specific heads to compensate the potential performance degradation caused by gradient conflicts.
To address the above limitations, we aim to increase the mutual information between the model’s input features and the task-specific outputs. Partitioning the parameter space into distinct subspaces has proven to be effective for enhancing mutual information, particularly when addressing diverse tasks (Ueda and Nishikawa, 2020). It reduces task interference and enables the model to learn task-specific predictions through different task optimization subspaces. To naturally achieve such partitioning, we design a generative paradigm that leverages instruction tuning in pre-trained generative models. Instruction tuning partitions the parameter space by introducing distinct prompt for each task, effectively tailoring a unique optimization subspace for S&R as illustrated in Figure 1(b).
To this end, we propose a novel Generative paradigm for unifying Search and Recommendation, abbreviated as GenSR, which leverages task-specific prompts to partition the model’s parameter space into subspaces, thereby enhancing mutual information (see detailed analysis in §5). To construct effective subspaces for each task, GenSR first prepares informative representations for each subspace and then optimizes both subspaces in one unified model. Specifically, GenSR consists of two main modules: (1) Dual representation learning: GenSR independently models historical dual representations that encompass sufficient information for both S&R tasks. Specifically, we simultaneously learn item representations from both collaborative filtering (CF) view and semantic view. Meanwhile, to capture user preference from extensive S&R historical interactions, a soft-filtering strategy is designed to prioritize the most relevant representations. (2) S&R task unifying: GenSR integrates previously learned representation with task-specific prompts and feeds them into a shared generative model to perform both tasks. Specifically, it employs contrastive learning to align the CF and semantic representations for the same users and uses instruction tuning to effectively generate task-specific outputs.
In summary, the main contributions are threefold: (1) We propose GenSR, a generative paradigm designed to unify S&R tasks through dual representation learning and S&R task unifying. (2) We introduce a novel perspective for comparing discriminative and generative paradigms based on mutual information. (3) We conduct comprehensive experiments on two public datasets, including S&R performance analysis, mutual information estimation, optimization space visualization, and ablation study of different modules, to validate GenSR’s effectiveness and efficiency compared to other baselines across both S&R tasks.
2. Related Work
Unifying search and recommendation. Unifying S&R by leveraging data from both tasks can enhance user modeling and item understanding, broadly categorized into two approaches: 1) Search-enhanced recommendation (Wu et al., 2019), which uses user search behavior as supplementary to uses’ representation, thus assisting recommendation. SESRec (Si et al., 2023) utilizes contrastive learning to extract similar and dissimilar user interests from S&R, and then leverages users’ search interests for recommendation. 2) Unifying S&R (Yao et al., 2021), which shares S&R data and utilizes one model to boost the performance on both two tasks. For example, UniSAR (Shi et al., 2024) models fine-grained user behavior transitions between S&R through extraction, alignment, and fusion. However, they follow the discriminative paradigm, which introduces gradient conflict and manual design complexity. Recently, Penha et al. (2024) proposed a generative model for unifying S&R. While promising, it treats S&R data independently, ignoring their joint interaction history—a rich source of contextual signals for user modeling.
In contrast, we introduce a generative paradigm that enhances mutual information and jointly leverages S&R interactions to better model user and item representations.
Generative recommendation. Pre-trained generative models have garnered attention in recommendation for their contextual understanding and global knowledge (Lin et al., 2024). Early works like P5 (Geng et al., 2022), TALLRec (Bao et al., 2023b), and M6-Rec (Cui et al., 2022) fine-tune these models with prompts and recommendation data. Later methods such as BIGRec (Bao et al., 2023a), IDGenRec (Tan et al., 2024), and TIGER (Rajput et al., 2023) enhance grounding in item space and semantic relevance. Recent efforts further highlight the critical role of CF information in generative recommendation, as CF plays a crucial role in understanding user and item representations in recommendation systems (Wang et al., 2024; Kim et al., 2024). However, integrating search data into generative recommendations remains underexplored and holds potential for capturing users’ immediate intent to boost personalization.
Personalized search. Personalized search aims to deliver search results tailored to users’ preferences, behaviors, and contextual information. For instance, HEM learns user representations from user profiles to provide personalized search rankings (Ai et al., 2017). TEM enhances this by incorporating user search histories, allowing it to better capture users’ preferences and historical behaviors (Bi et al., 2020).
Generative model for ranking. Due to the advanced contextual understanding capabilities of pre-trained language models, generative models for ranking have emerged as a novel approach to directly model the relevance between queries and items by framing ranking as a sequence generation task. For example,  Nogueira et al. (2020) adapts T5 for document ranking by training it to generate relevance labels (“true” or “false”) for query-document pairs. Similarly,  Ferraretto et al. (2023) introduces a generative ranking model trained to output both relevance labels and explanations for query-document pairs. Despite these advances, it’s promising to incorpoate recommendation data to extract users’ preferences which helps to boost the performance on the search side.
3. Task Formulation
We formulate the task of unifying S&R under both discriminative and generative paradigms. The objective is to integrate users’ historical S&R interactions into a single model capable of handling both tasks based on user needs. Specifically, let $\mathcal{U}$ and $\mathcal{I}$ be the sets of users and items. For a user $u\in\mathcal{U}$, the model predicts preference for a candidate item $i_{c}$ in either a search or recommendation context. To achieve this, the model takes as input three types of information for each user $u$: (1) Interaction history $H_{u}=\{(i_{1},b_{1}),(i_{2},b_{2}),\cdots,(i_{N},b_{N})\}$, where $i_{n}\in\mathcal{I}$ is the $n-$th item in the interaction history, $N$ is the total number of interactions, $b_{n}$ is the type of behavior for $n-$th item, with ”src” indicating search and ”rec” indicating recommendation. (2) Candidate item $i_{c}$, which is the item candidate pool used for evaluating whether the user $u$ will like (i.e., click) this item. (3) Query $q$, issued if the current task is search; otherwise, $q$ is empty for recommendation.
Discriminative paradigm. This paradigm encodes inputs via shared encoders and applies task-specific heads. The model minimizes the combined cross-entropy loss for both tasks:
(1) $\mathop{\min}_{\Phi}\{\mathcal{L}_{\Phi}=\mathcal{L}{s}+\gamma\mathcal{L}{r}\},$
where $\Phi$ denotes model parameters, $\gamma$ controls the trade-off between S&R tasks, and $\mathcal{L}_{s}$, $\mathcal{L}_{r}$ are cross-entropy losses for S&R.
Generative paradigm. Differently, in the generative paradigm, the input information is transformed into a unified structured prompt $x$, which is tokenized before being fed into the model for instruction tuning. The model generates “Yes” or “No” to indicate user preference via constrained generation.
The training objective minimizes the negative the negative log-likelihood of the users’ preference $y$ conditioned on input $x$:
(2) $\mathop{\min}_{\Theta}\{\mathcal{L}_{\Theta}=-\sum_{t=1}^{|y|}\log P_{\Theta}(y_{t}|y_{<t},x)\},$
where $\Theta$ denotes model parameters, $x$ and $y$ denote the “Instruction Input” and “Instruction Output” in the self-instruct data for instruction tuning, $y_{t}$ is the $t$-th token of $y$, and $y_{<t}$ represents all tokens preceding $y_{t}$.
4. Method
Figure 2. GenSR structure. Given S&R history, GenSR uses dual representation learning to obtain CF and semantic embeddings, and unifies both tasks via contrastive learning combined with instruction tuning to generate task-specific outputs.
We propose GenSR, a novel generative paradigm for unifying search and recommendation (S&R). GenSR has two modules: 1) Dual Representation learning, which independently learns item representations from both the CF view and the semantic view. 2) S&R Task Unifying, which utilizes contrastive learning together with instruction tuning to generate task-specific outputs effectively. The overall architecture of GenSR is shown in Figure 2.
4.1. Dual representation learning
To derive informative dual representation for each subspace, GenSR first models historical dual-view representation to encompass sufficient information for both S&R tasks, and then employs soft-filtering strategy to capture use preference from extensive S&R historical representations.
4.1.1. Dual-view representation
When unifying S&R tasks, it is crucial to model item representations that encapsulate sufficient information to effectively support both tasks. A key consideration in this process is the distinct focus of S&R tasks on item representations. Specifically, for the search task, item representations rely heavily on semantic matching between queries and items, as users actively search for specific items based on their explicit information needs. In contrast, for the recommendation task, item representations are mainly shaped by CF information derived from historical user interactions (Zhao et al., 2023), as recommendations are passively received by users who often lack explicit information needs.
This distinction underscores the need to integrate both semantic and CF information to construct comprehensive item representations. To tackle this challenge, we propose a dual-view representation method that separately captures semantic and CF information from two views.
CF view. To model the CF view for the dual-view representation, we begin by extracting the S&R history $H_{u}=\{i_{1},i_{2},\cdots,i_{N}\}$ for a given user $u$. This history represents all chronologically ordered S&R interactions, where $N$ denotes the total number of interactions recorded in the user’s history $H_{u}$.
From this history, we compute the corresponding CF representations $E_{u}^{\text{cf}}=\{e_{1},e_{2},\cdots,e_{N}\}^{T}\in\mathbb{R}^{N\times d}$ using a pre-trained CF embedding table ||×$\mathcal{E}\in\mathbb{R}^{|\mathcal{I}|\times d}$. Here, each $e_{i}$ is a $d$-dimensional embedding, $|\mathcal{I}|$ denotes the total number of items in the corpus, and $d$ represents the dimensionality of the CF embeddings. [^1] The CF embedding table $\mathcal{E}$ is derived from training LightGCN using only the recommendation behaviors in the dataset (He et al., 2020).
To capture the sequential dependencies within the S&R history, the representations $E_{u}^{\text{cf}}$ are further processed using a transformer encoder. The encoder comprises a Multi-Head Self-Attention (MSA) layer and a Feed-Forward Network (FFN). Specifically, the MSA layer takes $E_{u}^{\text{cf}}$ as the input, where the query ($Q$), key ($K$), and value ($V$) matrices are all set to $E_{u}^{\text{cf}}$. Then the CF representation $C_{u}\in\mathbb{R}^{N\times d}$ is computed as:
(3) $C_{u}=FFN(MSA(E_{u}^{\text{cf}},E_{u}^{\text{cf}},E_{u}^{\text{cf}})),$
Semantic view. To model the semantic view for the dual-view representation, we process the user’s S&R history $H_{u}=\{i_{1},i_{2},\cdots,i_{N}\}$, which contains all chronologically ordered interactions, where $N$ is the total number of interactions. Semantic representations $E_{u}^{\text{sem}}=\{s_{1},s_{2},\cdots,s_{N}\}^{T}\in\mathbb{R}^{N\times d^{\prime}}$ are derived by tokenizing the textual descriptions of items using a generative model tokenizer. Here, $d^{\prime}$ is the embedding dimensionality of the generative model, and each $s_{i}$ is a $d^{\prime}$-dimensional embedding capturing the item’s semantic information.
To capture the sequential dependencies, the representations $E_{u}^{\text{sem}}$ are fed into a transformer encoder. The transformer encoder consists of a Multi-head Self-Attention (MSA) layer and a Feed-Forward layer (FFN). Specifically, MSA takes $E_{u}^{\text{sem}}$ as the input, where $Q=K=V=E_{u}^{\text{sem}}$. Then the semantic representation $S_{u}\in\mathbb{R}^{N\times d^{\prime}}$ is computed as:
(4) $S_{u}=FFN(MSA(E_{u}^{\text{sem}},E_{u}^{\text{sem}},E_{u}^{\text{sem}})).$
4.1.2. S&R history soft filtering
When combining S&R behaviors together in the user history sequence, it is essential to effectively capture user preferences from an extensive history of S&R interactions. The fusion of these two tasks introduces a larger volume and greater complexity of user behaviors, making it crucial to extract meaningful patterns from the multi-task interaction history. To address this challenge, we propose a history soft filtering strategy, which selectively soft-filters relevant items from the extensive S&R interaction history based on the user’s specific information needs.
Intuitively, after obtaining the learned CF representation $C_{u}$ and semantic representation $S_{u}$ from the dual-view representation, we calculate the importance score of each historically interacted item with respect to the candidate item $i_{c}$. Specifically, for the CF side, given a user $u$, the CF representation $C_{u}=\{h_{1},h_{2},\dots,h_{N}\}$ contains the encoded CF embeddings of $N$ historically interacted items, where $h_{j}$ is the encoded CF embedding of the $j$-th item. To determine the relevance of each $h_{j}$ to the candidate item $i_{c}$ with CF embedding $e_{i_{c}}$, we compute the importance score $\alpha_{j}$ using a multi-head cross-attention (MCA) module:
(5) $\alpha_{j}=\exp\left(\text{MCA}(e_{i_{c}},h_{j})\right)/\sum_{k=1}^{N}\exp\left(\text{MCA}(e_{i_{c}},h_{k})\right),$
where $\alpha_{j}$ is the importance score for the $j$-th historical item, and $\text{MCA}(e_{i_{c}},h_{j})$ represents the attention score between the candidate item’s CF embedding $e_{i_{c}}$ and the $j$-th item CF embedding $h_{j}$.
We then apply soft-filtering to the S&R interaction history using the calculated importance scores $\alpha_{j}$:
(6) $C_{u}^{\text{filter}}=\sum_{j=1}^{N}\alpha_{j}\cdot W_{V}h_{j},$
where $W_{V}$ is a learned projection matrix associated with $C_{u}$, and $h_{j}$ is the $j$-th optimized CF embedding from $C_{u}$.
Similarly, we perform the same soft-filtering for the semantic side, obtaining $S_{u}^{\text{filter}}$. By learning both the filtered CF and semantic representations of historical S&R interactions, we obtain relevant dual representation that supports both S&R tasks.
4.2. S&R task unifying
After preparing the dual representation, GenSR integrates these learned representation with task-specific prompts, which are then fed into a shared generative model to unify S&R tasks.
4.2.1. Training
We use instruction tuning to train the generative model for performing S&R tasks, enabling the parameter space to naturally partition into subspaces optimized for different tasks. Specifically, after obtaining $C_{u}$ and $S_{u}$ from the dual representation learning, we construct a task-specfic prompt that incorporates these representations, along with the query $q$ (for search tasks only), the candidate item’s CF embeddings $e_{i_{c}}$ and semantic embeddings $s_{i_{c}}$. Formally, we use generative model tokenizer to tokenize the textual content of the prompt, which contains textual instructions and the query (if applicable). The result of the tokenization is represented as: $P_{t}=[p_{1}^{t},\langle PH1\rangle,p_{2}^{t},\langle PH2\rangle,p_{3}^{t},\langle PH3\rangle,p_{4}^{t},\langle PH4\rangle,p_{5}^{t}]$, where task $t\in\{S,R\}$, $p_{k}^{t}$ denotes $k$-th segments of prompt text tokens for task $t$, and $\langle PH\rangle$ represents placeholders that will be replaced with other feature embeddings later. [^2] The five segments of prompt text tokens ($p_{1}^{t}$ to $p_{5}^{t}$) are used because, in our setting, we only have four feature embeddings to insert into the prompt. Therefore, the prompt is divided into five segments to interleave the embeddings properly. We then encode the prompt into a sequence of embeddings and replace the placeholders with learned representation and candidate item embeddings for each task:
(7) $E_{t}=[e_{{1}}^{t},C_{u},e_{{2}}^{t},S_{u},e_{{3}}^{t},e_{i_{c}},e_{{4}}^{t},s_{i_{c}},e_{{5}}^{t}],\quad\text{where}\quad t\in\{S,R\}.$
Here, $e_{k}^{t}\in\mathbb{R}^{1\times d_{2}}$ represents the embedding for token $p_{1}^{k}$, obtained via the embedding lookup in the generative model. The final tokenized embedding $E_{t}$ is fed into the generative model as input for instruction tuning. This approach allows the model to adaptively leverage both CF and semantic information, while employing task-specific prompts to partition the model’s parameter space into subspaces for optimizing different tasks.
The overall loss function for the GenSR model consists of two components: 1) a task-specific generation loss for S&R, $\mathcal{L}_{\text{S\&R}}$, and 2) a contrastive learning loss, $\mathcal{L}_{c}$, which aligns the CF and semantic representations. The combined loss is:
(8) $\mathcal{L}=\mathcal{L}_{\text{S\&R}}+\beta\mathcal{L}_{c},\text{where}$
(9) $\mathcal{L}_{\text{S\&R}}=-\sum_{t=1}^{|y|}\log P_{\Theta}(y_{t}|y_{<t},E)\},$
where $\beta$ is the loss balance factor that controls the weight of the contrastive loss relative to the S&R task loss, and $|y|$ is the length of the target sequence, $y_{t}$ is the token at position $t$ in the target sequence, and $y_{<t}$ represents all tokens preceding $t$. The conditional probability $P_{\Theta}(y_{t}\mid y_{<t},E)$ is computed by the generative model, parameterized by overall model parameters $\Theta$, given the input tokenized embedding $E$. This term encourages the model to maximize the likelihood of generating the correct output sequence for S&R.
4.2.2. Contrastive learning between CF and Semantic views
Since we design dual representation learning from the CF and semantic views, it is important to notice that these perspectives provide complementary insights into user preferences. Aligning these two views enables the model to integrate their respective strengths, resulting in comprehensive and robust user representations that enhance generalization across diverse S&R tasks.
To effectively align user representations from both the CF and semantic views, we employ contrastive learning. This approach encourages the CF and semantic representations of the same user $u$ to be closer in the embedding space, while ensuring that the CF or semantic representations of different users $u^{\prime}$ (within the same batch) are pushed further apart.
Specifically, for a given user $u$, with CF representation $C_{u}$ and semantic representation $S_{u}$ obtained from dual-view representation, the goal is twofold: 1) Minimize the distance between $C_{u}$ and $S_{u}$ (positive pair), ensuring that the CF and semantic embeddings of the same user are closely aligned, effectively transferring complementary information between views. 2) Maximize the distance between $C_{u}$ (or $S_{u}$) and the representations $C_{u^{\prime}}$ or $S_{u{\prime}}$ of other users $u^{\prime}$ in the batch (negative pairs). This enforces distinctiveness between users, preventing overlapping embeddings that could reduce model performance in S&R tasks. The contrastive loss for this alignment is calculated as follows:
[…]
where $C_{u}$ and $S_{u}$ denote the CF and semantic embeddings for the same user $u$, while $u^{\prime}$ refers to other users in the batch $B$. The function $\text{sim}(\cdot,\cdot)$ is implemented as cosine similarity, measuring the similarity between two embeddings. $\tau$ is a temperature factor that adjusts the sharpness of the similarity distribution, controlling how much emphasis is placed on high-similarity pairs.
5. Mutual Information Analysis of GenSR
After illustrating the detailed methodology of GenSR, we conduct an in-depth analysis of mutual information to support our claim that the proposed generative paradigm effectively enhances mutual information compared to traditional discriminative methods. We begin by defining mutual information from an information-theoretic perspective, and then present both theoretical and empirical evidence to substantiate this claim in the following sections.
5.1. Information-theoretic perspective
From an information-theoretic standpoint, the representational capability of a unified S&R model can be examined via the mutual information between the input $X$ and the output $Y_{t}$ for a task $t\in\{S,R\}$. Mutual information, denoted as $I(X;Y_{t})$, measures how much information the input $X$ provides about the output $Y_{t}$, quantifying their dependence. A higher mutual information indicates that the input $X$ contains more predictive information about the output $Y_{t}$, reflecting a stronger alignment between inputs and task-specific outputs, i.e., higher model representation capability. Conversely, lower mutual information implies that predicted output fail to fully capture the distribution and pattern of input, reflecting a lower model representation capability. Therefore, mutual information $I(X;Y_{t})$ serves as a good indicator for model representation capability (Bachman et al., 2019).
5.2. Theoretical analysis
In this section, we aim to show the effectiveness of our model GenSR, based on the aforementioned mutual information. Specifically, we conduct a theoretical analysis to show this utilizing the following lemma:
Lemma 5.1.
Let the input $X\sim\mathcal{N}(\mu,\Sigma)$, and assume that noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}I)$ is independent Gaussian noise. Consider two paradigms: generative (with parameters $\phi$) and discriminative (with parameter $\theta$). The mutual information between the input and output satisfies:
(11) $I_{\theta}(X;Y_{S})+I_{\theta}(X;Y_{R})\leq I_{\phi}(X;Y_{S})+I_{\phi}(X;Y_{R}).$
Here, $I_{\phi}(X;Y_{S})$ and $I_{\phi}(X;Y_{R})$ denote the mutual information between input and output under the generative paradigm for S&R tasks, respectively, while $I_{\theta}(X;Y_{S})$ and $I_{\theta}(X;Y_{R})$ denote mutual information under the discriminative paradigm.
We provide a detailed proof of Lemma 5.1 from the view of parameter approximation separability in the Appendix. Intuitively, we prove that generative paradigm can enhance mutual information in S&R scenario compared with discriminative paradigm under certain input distribution assumption.
5.3. Empirical analysis
Table 1. Mutual Information Estimation.
Amazon KuaiSAR
Method Search Rec. Search Rec.
UniSAR 0.9003 0.8739 0.8514 0.8475
GenSR 0.9378 0.9217 0.9251 0.9168
To further support our theoretical claim and verify the conclusion in a broader complex real-world scenario, we employ Mutual Information Neural Estimation (MINE) to empirically estimate the practical mutual information between the input $X$ and the task-specific outputs $Y_{S}$ and $Y_{R}$ (Belghazi et al., 2018). Our empirical findings are consistent with the theoretical conclusion. Specifically, we use the Donsker-Varadhan representation of KL divergence to estimate the lower bound of mutual information as:
(12) $I(X;Y_{t})\geq\mathbb{E}_{P(X,Y_{t})}[f(x,y_{t})]-\log\mathbb{E}_{P(X)P(Y_{t})}[e^{f(x,y_{t})}],$
where $f(x,y_{t})$ is a neural network function trained to maximize this bound.
In our experiments, we calculate mutual information for both S&R tasks under the discriminative and generative paradigms. For each task, we estimate $I(X;Y_{t})$ by sampling data from the joint distribution $P(X,Y_{t})$ and comparing it with samples from the independent marginal distributions $P(X)$ and $P(Y_{t})$. This process provides a quantitative measure of how well the input and output distributions are aligned in each paradigm. We compare GenSR with the most powerful discriminative paradigm baseline UniSAR in two public datasets: Amazon and KuaiSAR (see detailed dataset description and baseline description in Section 6.1). The results, presented in Table 1, show that the generative paradigm consistently achieves higher mutual information for both tasks. This improvement aligns with our theoretical analysis, which is attributed to the task-specific subspaces.
6. Experiments
In this section, we conduct a comprehensive experimental study to address the following research questions. RQ1: How does the performance of GenSR compare to other baseline methods across various datasets in both recommendation and search scenarios? RQ2: How well can GenSR enhance mutual information compared to discriminative approaches? RQ3: What is the impact of different strategies (e.g., dual-view representation and S&R history soft filtering) within GenSR on its overall performance?
Table 2. Overall performance of GenSR and other baselines under recommendation scenario. The best and second-best results are highlighted in bold and underlined fonts, respectively. * denotes statistically significant improvements (t-test, $p$ 0.01).
Recommendation Joint Search and Recommendation
Dataset Metric DIN GRU4Rec SASRec BERT4Rec FMLP CoLLM IDGenRec NRHUB SeqRec SESRec JSR USER UnifiedSSR UniSAR BSR GenSR
Amazon R@5 0.5170 0.4949 0.5295 0.5311 0.5356 0.5800 0.2305 0.4988 0.5401 0.5623 0.5467 0.5441 0.5196 0.5874 0.1417 0.6102*
R@10 0.6525 0.6548 0.6772 0.6658 0.6879 0.6811 0.2818 0.6503 0.6860 0.6864 0.6779 0.6854 0.6707 0.7020 0.2278 0.7200*
N@5 0.3726 0.3388 0.3747 0.3954 0.3739 0.4651 0.1791 0.3487 0.3859 0.4245 0.3970 0.3964 0.3662 0.4513 0.0899 0.4922*
N@10 0.4165 0.3907 0.4225 0.4390 0.4232 0.5001 0.1955 0.3977 0.4333 0.4648 0.4396 0.4422 0.4151 0.4885 0.1177 0.5278*
KuaiSAR R@5 0.4509 0.3764 0.4065 0.3699 0.4292 0.4092 0.3862 0.3920 0.4956 0.4791 0.4086 0.3981 0.5169 0.2180 0.5235*
R@10 0.6179 0.5788 0.6007 0.5885 0.6159 0.5915 0.5610 0.5890 0.6643 0.6453 0.5627 0.5939 0.6792 0.3395 0.6819*
N@5 0.3104 0.2345 0.2671 0.2381 0.2851 0.3199 0.2572 0.2552 0.3432 0.3315 0.2820 0.2617 0.3632 0.1472 0.3686*
N@10 0.3643 0.3087 0.3298 0.3083 0.3453 0.3683 0.3136 0.3186 0.3978 0.3853 0.3318 0.3249 0.4158 0.1860 0.4199*
Table 3. Overall performance of GenSR and other baselines under search scenario. Bold indicates the best result and underline the second-best, with * denoting statistically significant improvements (t-test, $p$ 0.01).
Search Joint Search and Recommendation
Dataset Metric QEM HEM AEM ZAM TEM CoPPS JSR USER UnifiedSSR UniSAR BSR GenSR
Amazon R@5 0.7100 0.6778 0.7095 0.7109 0.8185 0.8169 0.7038 0.7631 0.7744 0.8190 0.4786 0.9782*
R@10 0.8186 0.8267 0.8443 0.8468 0.9051 0.9051 0.8225 0.8697 0.8812 0.8977 0.6015 0.9664*
N@5 0.5066 0.4736 0.5114 0.5147 0.6303 0.6281 0.5173 0.6000 0.5847 0.6875 0.3485 0.9853*
N@10 0.5422 0.5221 0.5554 0.5590 0.6587 0.6570 0.5563 0.6348 0.6196 0.7132 0.3883 0.9687*
KuaiSAR R@5 0.6020 0.6505 0.5956 0.6117 0.6502 0.6616 0.7162 0.7304 0.7377 0.7476 0.1573 0.7661*
R@10 0.7182 0.7653 0.7182 0.7344 0.7632 0.7707 0.7961 0.8149 0.8320 0.8369 0.2075 0.8372
N@5 0.4575 0.5029 0.4415 0.4560 0.4887 0.4977 0.5962 0.6069 0.5991 0.6417 0.1359 0.7564*
N@10 0.4953 0.5400 0.4812 0.4959 0.5254 0.5331 0.6221 0.6342 0.6297 0.6708 0.1520 0.7655*
6.1. Experimental setups
6.1.1. Datasets.
We evaluate GenSR on the two publicly accessible datasets that contain both S&R histories. 1) Amazon [^3] https://jmcauley.ucsd.edu/data/amazon/. is a recommendation review dataset from Amazon. We use the ”Kindle Store” 5-core sub-dataset from the dataset, which contains user reviews and ratings for Kindle books. Adopting the procedure from prior work (Si et al., 2023; Shi et al., 2024), we first generate synthetic search behaviors and queries based on this recommendation dataset and then apply a leave-one-out strategy for dataset splitting. The detailed process for generating search behaviors is described in (Ai et al., 2017). This automatically constructed search dataset has been widely adopted by the product search community and prior work on unifying search and recommendation (Si et al., 2023; Shi et al., 2024; Ai et al., 2017; Van Gysel et al., 2016; Guo et al., 2019). It contains 68,223 users, 61,934 items, 4,298 queries, 989,618 recommendation interactions, and 934,664 search interactions. 2) KuaiSAR [^4] https://kuaisar.github.io/. is a micro-video dataset that offers authentic S&R behaviors of users, featuring detailed item categories and real user queries. We preprocess the data to include only users showing both S&R behaviors. Following (Shi et al., 2024), we filter for users and items with at least five interactions and allocate the latest day’s data for testing, the previous day for validation, and the earlier data for training. It contains 25,877 users, 6,890,707 items, 453,667 queries, 14,605,716 recommendation interactions, and 5,059,169 search interactions.
6.1.2. Baselines.
We compare our method with the state-of-the-art recommendation, search, and unifying methods.
The recommendation baselines are: 1) DIN (Zhou et al., 2018) generate ad-specific user interest representations from historical behaviors. 2) GRU4Rec (Hidasi et al., 2016) utilizes RNNs for session-based recommendations. 3) SASRec (Kang and McAuley, 2018) applies self-attention to model long-term user behavior. 4) BERT4Rec (Sun et al., 2019) uses bidirectional modeling to predict masked items in sequences. 5) FMLP (Zhou et al., 2022) uses an all-MLP model that applies frequency-domain filtering to sequence recommendation. 6) CoLLM (Zhang et al., 2023) integrates collaborative embeddings into LLMs for recommendation. 7) IDGenRec (Tan et al., 2024) generates personalized recommendations using textual item IDs. [^5] We only implement IDGenRec on Amazon due to its requirement for explicit semantic textual information of items, which is unavailable for KuaiSAR.
The search baselines are: 8) QEM (Ai et al., 2019) assesses item-query matching scores. 9) HEM (Ai et al., 2017) utilizes a hierarchical embedding model that captures the semantics of queries, items, and users. 10) AEM (Ai et al., 2019) offers personalizes search results through an attention mechanism based on query and user data. 11) ZAM (Ai et al., 2019) enhances AEM by appending a zero vector for better item differentiation. 12) TEM (Bi et al., 2020) implements a Transformer-based embedding model for personalized product search. 13) CoPPS (Dai et al., 2023) introduces contrastive learning to refine search outcomes from sparse and noisy data.
The unifying baselines include two categories: (a) Search enhanced recommendation: 14) NRHUB (Wu et al., 2019) enhances user models with search queries for improved recommendation. 15) SeqRec (He et al., 2022) uses queries to drive user behavior modeling with transformers. 16) SESRec (Si et al., 2023) uses contrastive learning to sort user interests from S&R data. (b) Jointly unified S&R: 17) JSR (Zamani and Croft, 2018) jointly models S&R tasks with a shared neural network. 18) USER (Yao et al., 2021) merges user behaviors from both domains into a cohesive sequence for joint modeling. 19) UnifiedSSR (Xie et al., 2024) employs an intent-oriented session modeling module to infer user intent. 20) UniSAR (Shi et al., 2024) models user behavior transitions between S&R. 21) BSR (Penha et al., 2024) employs atomic item IDs to do the S&R task in generative model.
6.1.3. Experimental settings.
Following previous work (Shi et al., 2024; Xie et al., 2024; Yao et al., 2021), we randomly sample 99 negative items that the user has not interacted with and combine them with the ground truth item to create the candidate item list. All the models are then tasked with ranking these candidate items to evaluate top-K S&R performance using Recall@K and NDCG@K with $K=\{5,10\}$. To ensure a fair comparison, we implement constrained beam search during inference for two full-ranking generative baselines, IDGenRec and GenSR, which guarantees that the generated results are restricted to the valid candidate set. Regarding our proposed GenSR, we select flan-t5-base as the backbone. The best hyper-parameters are selected with the searching scopes as follows: the temperature factor $\tau$ and the loss balance factor $\beta$ are tuned within $\{0.01,0.03,0.05,0.07,0.09\}$ and $\{0.001,0.01,0.1,0.15,0.2,0.25,0.3\}$.
6.2. Overall performance (RQ1)
We conduct experiments in both S&R scenarios. The results, illustrated in Table 2 and Table 3, yield several key observations:
  • GenSR consistently outperforms all baselines across all datasets in both S&R tasks. This superior performance can be attributed to GenSR’s dual representation learning and S&R task unifying under the superior generative paradigm.
  • Unified methods for S&R generally outperform baselines that focus on only one task. This highlights that leveraging shared user behaviors and item interactions in the S&R scenario can mutually enhance performance in both tasks.
  • The performance of IDGenRec is relatively weaker, potentially due to its reliance on semantic information for the recommendation task, which may limit its ability to capture collaborative filtering information explicitly. Similarly, BSR’s performance may be constrained by its use of atomic item IDs, which lack semantic information to enrich item representations further.
6.3. In-depth analysis
Figure 3. Parameter space visualization under KuaiSAR.
6.3.1. Parameter space visualization analysis (RQ2)
We analyze the parameter space to support three insights: (1) S&R tasks exhibit distinct data distributions; (2) discriminative paradigm forces both tasks into a shared space, causing gradient conflicts; and (3) generative paradigm naturally separates this space, reducing such conflicts. Experiments are based on KuaiSAR, the only real-world public dataset. To examine the raw data distributions of S&R tasks, we utilize Kernel Density Estimation (KDE) to approximate $P(X_{S},Y_{S})$ and P$(X_{R},Y_{R})$ (Terrell and Scott, 1992), using concatenated $X$ and $Y$ vectors. After dimensionality reduction via t-SNE(Van der Maaten and Hinton, 2008), we compute KDE in the embedded space. The resulting contours (Figure 3a) reveal distinct patterns, confirming distributional differences of two tasks. Given the observed differences in task distributions, we then examine parameter space overlap under the discriminative setting by projecting task-specific hidden states using PCA (Figure 3b). The significant overlap reflects optimization conflicts due to mismatched task distributions. In contrast, the generative paradigm naturally partitions the parameter space into task-specific subspaces that align with the separate distributions of S&R tasks (Figure 3c). This separation enables independent optimization for each task, effectively mitigating task conflicts and enhancing the model’s ability to handle both tasks within a unified framework.
Figure 4. Ablation study of different components in GenSR.
6.3.2. Ablation study (RQ3)
We conduct an ablation study to verify the effectiveness of each part of GenSR. We assess the unique contributions of dual-view representation, S&R history soft filtering, and contrastive learning. The results, presented in Figure 4, reveal the following findings: 1) Preserving only one representation view reduces performance on both datasets. This result underscores the advantage of independently learning item representations from the CF and semantic views, enabling the model to effectively integrate the diverse information needed to support both S&R tasks. 2) Excluding S&R soft filtering also leads to a decline in performance, which demonstrates the effectiveness of prioritizing relevant information from interaction histories. 3) Removing contrastive learning similarly decreases performance, which highlights the importance of aligning the CF-view and semantic view for the same user.
Figure 5. Hyper-parameters analysis under Amazon.
6.3.3. Hyper-parameter analysis (RQ3)
We examine the impact of two hyper-parameters of GenSR in Figure 5 using the Amazon dataset (excluding KuaiSAR due to similar trends): the temperature factor $\tau$ used in contrastive learning and the loss balance factor $\beta$ controlling the weight of the contrastive loss relative to the S&R task loss. The result shows that: 1) Varying $\tau$ affects how sharply similarities are emphasized among CF and semantic embeddings. Smaller $\tau$ values highlight high-similarity pairs more strongly but risk overfitting, while larger $\tau$ values foster more even attention at the expense of less pronounced discrimination. 2) Similarly, $\beta$ determines the trade-off between focusing on embedding alignment versus optimizing the main S&R tasks. An excessively high $\beta$ could overemphasize embedding alignment, marginalizing S&R performance; a too-low $\beta$ may fail to leverage the complementary strengths of the two views.
7. Conclusion
In this work, we have introduced GenSR, a novel generative paradigm designed to unify search and recommendation tasks, which leverages task-specific prompts to partition the model’s parameter space into subspaces, thereby enhancing mutual information. To construct effective subspaces for each task, GenSR first prepares informative representations for each subspace through dual representation learning and then optimizes both subspaces in one unified model. Comprehensive experiments on two public datasets have validated GenSR’s effectiveness, showing consistent improvements over state-of-the-art methods across S&R tasks.
While our framework is grounded in an information-theoretical analysis and employs task-specific subspaces to increase mutual information, there is potential to further optimize these subspaces from an optimization perspective. In future work, we aim to explore advanced optimization techniques to improve the alignment and mutual information between these task-specific subspaces.
Appendix
Proof of Lemma 5.1
We consider two paradigms: the discriminative paradigm with parameter $w_{\theta}$ and the generative paradigm with parameter $w_{\phi}$. Let the true parameters be $w_{\theta}^{*}$ and $w_{\phi}^{*}$, and let their estimates from data be $\hat{w}_{\theta}$ and $\hat{w}_{\phi}$ respectively.
The parameter approximation is obtained by minimizing their respective loss functions: [^6] For binary generation tasks (e.g., “yes” vs. “no”), the negative log-likelihood in the generative setting aligns with the binary cross-entropy loss of the discriminative setting, allowing the same format of loss.
$\hat{w}_{\theta}=\arg\min_{w_{\theta}}L(w_{\theta}),\quad\hat{w}_{\phi}=\arg\min_{w_{\phi}}L(w_{\phi}).$
In the binary classification setting, it has been proved that the discriminative parameter can be represented as the difference of the generative parameters across classes (i.e., $w_{\theta}^{*}=w_{\phi,1}^{*}-w_{\phi,0}^{*}$, where $w_{\phi,0}^{*}$ and $w_{\phi,1}^{*}$ denote the true parameters for class 0 and class 1, ) (Prasad et al., 2017). Hence, comparing parameter approximation errors for the two paradigms is meaningful. To do that, we expand the loss functions around the respective true parameters using Taylor’s expansion. We define the approximation errors as follows: $\Delta_{D}=\hat{w}_{\theta}-w_{\theta}^{*}$, $\Delta_{G}=\hat{w}_{\phi}-w_{\phi}^{*}$. For each paradigm, we have:
$\nabla L_{n}(w_{\theta}^{*})+\nabla^{2}L_{n}(w_{\theta}^{*})\Delta_{D}+R_{D}(\Delta_{D})=0,$
$\nabla L_{n}(w_{\phi}^{*})+\nabla^{2}L_{n}(w_{\phi}^{*})\Delta_{G}+R_{G}(\Delta_{G})=0,$
where $R_{D}(\Delta_{D})$ and $R_{G}(\Delta_{G})$ are the corresponding higher-order remainder terms that capture the nonlinear and cross-dimensional interactions. Following (Prasad et al., 2017), we assume that the remainder terms satisfy a uniform bound in the local neighborhood: for any fixed small perturbation $\Delta$, then:
$\|R_{G}(\Delta)\|_{1}\leq\frac{1}{\gamma_{G}}\|\Delta\|_{1}^{\alpha}\quad\text{and}\quad\|R_{D}(\Delta)\|_{1}\leq\frac{1}{\gamma_{D}}\|\Delta\|_{1}^{\alpha},\quad\alpha\geq 2,$
where $\gamma_{G}$ and $\gamma_{D}$ are the separability constants for the generative and discriminative paradigms respectively. Since it has been proven that the generative paradigm enjoys higher separability (i.e. $\gamma_{G}>\gamma_{D}$(Prasad et al., 2017), we have $\|R_{G}(\Delta)\|_{1}<\|R_{D}(\Delta)\|_{1}$, for any fixed $\Delta$.
Assuming that the perturbations are small enough for the linear part of the expansion to dominate, we can apply the Taylor expansions:
$\Delta_{G}\approx-\left[\nabla^{2}L_{n}(w_{\phi}^{*})\right]^{-1}R_{G}(\Delta_{G}),\quad\Delta_{D}\approx-\left[\nabla^{2}L_{n}(w_{\theta}^{*})\right]^{-1}R_{D}(\Delta_{D}).$
Moreover, since the generative paradigm benefits from both explicit partitioning by prompts and higher separability, its Hessian takes a block-diagonal form and is better conditioned: [^7] Although this uniform bound is stated for an arbitrary (fixed) $\Delta$, in practice the error analysis is applied to the respective perturbations $\Delta_{D}$ and $\Delta_{G}$ in their own local neighborhood. Since both $\Delta_{D}$ and $\Delta_{G}$ are assumed to be small (and we are only using the bound locally), it is reasonable to apply the same type of bound separately to each—even though they are different. The uniformity of the bound assures us that the same structure holds regardless of which (small) perturbation is under consideration.
$\left\|\left[\nabla^{2}L_{n}(w_{\phi}^{*})\right]^{-1}\right\|\leq\left\|\left[\nabla^{2}L_{n}(w_{\theta}^{*})\right]^{-1}\right\|.$
Thus, combining these observations we obtain
$\|\Delta_{G}\|$ $=\left\|\left[\nabla^{2}L_{n}(w_{\phi}^{*})\right]^{-1}R_{G}(\Delta_{G})\right\|\leq\left\|\left[\nabla^{2}L_{n}(w_{\phi}^{*})\right]^{-1}\right\|\,\|R_{G}(\Delta_{G})\|_{1}$
$<\left\|\left[\nabla^{2}L_{n}(w_{\theta}^{*})\right]^{-1}\right\|\,\|R_{D}(\Delta_{D})\|_{1}\approx\|\Delta_{D}\|,$
where the last approximate equality holds when the remainder terms are small compared to the linear term. This chain of inequalities leads to the conclusion:
(13) $\|\hat{w}_{\phi}-w_{\phi}^{*}\|\leq\|\hat{w}_{\theta}-w_{\theta}^{*}\|.$
Consider a Gaussian model given by: $Y=w^{*}X+b^{*}+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ independent of $X$, $X\sim\mathcal{N}(\mu_{X},\sigma_{X}^{2})$. Building on prior work in information theory and statistical learning (e.g.,(Leung and Barron, 2006; Gabrié et al., 2018)), mutual information between input $X$ and prediction $\hat{Y}$, parameterized by model parameter $w$, is expressed as:
(14) $I(X;\hat{Y})=\frac{1}{2}\ln\left(1+\frac{w\sigma_{X}^{2}}{\sigma_{\epsilon}^{2}}\right).$
Suppose the noise variance $\sigma_{\epsilon}$ is the same in both paradigms (Hastie et al., 2009), as parameter estimates $\hat{w}$ approach the true parameter $w^{*}$, the mutual information $I(X;\hat{Y})$ approaches the theoretical upper bound $I(X;Y)$, which is identical for both paradigms. Given that we have established Eq.(13) and Eq.(14), the generative paradigm achieves lower approximation error, leading directly to higher mutual information:
$I_{w_{\theta}}(X;Y_{S})+I_{w_{\theta}}(X;Y_{R})\leq I_{w_{\phi}}(X;Y_{S})+I_{w_{\phi}}(X;Y_{R}).$